l. You loan from a bank the amount of P 100,000 with a rate of simple interest of 20% but the interest was deducted from the loan at the time the money was borrowed. If at the end of 1 year you have to pay the full amount of P 100 000 what is the actual rate of interest?
1
Expert's answer
2020-09-28T07:56:12-0400

Answer:

Simple interest on $100,000"=" $100,000 "\times" 20%"=" $20,000

Loan received at beginning of year"=" $100,000"-" $20,000 "=" $80,000

Actual interest rate (%) "=" [$(100,000"-" 80,000)"\/" $80,000] "\times" 100

"=" ($20,000"\/" $80,000) "\times" 100

"=" 25"%"%